OpenStudy (anonymous):

Which of these is equal to sin 21°? A) cos 21° B) cos 69° C) sin 69° D) tan 69°

OpenStudy (anonymous):

b

OpenStudy (anonymous):

B because, sinx = cos(90 - x) if x is in degrees sin21 = cos(90 - 21) = cos69

OpenStudy (freckles):

Recall sin(x)=cos(90-x)